1. Sarratt Corporation's contribution margin ratio is 77% and its fixed monthly expenses are $46,000. Assume that the company's sales for May are expected to be $105,000. Required: Estimate the company's net operating income for May, assuming that the fixed monthly expenses do not change.

2. In activity-based costing, the activity rate for an activity cost pool is computed by dividing the total overhead cost in the activity cost pool by:

  • the total direct labor-hours for the activity cost pool.

  • the direct labor-hours required by the product.

  • the machine-hours required by the product.

  • the total activity for the activity cost pool.

Which of the following is correct? The break-even point occurs on the CVP graph where:

  • total contribution margin equals total fixed expenses.

  • total profit equals total fixed expenses.

  • total profit equals total expenses.

  • total variable expenses equal total contribution margin.

Crich Corporation uses direct labor-hours in its predetermined overhead rate. At the beginning of the year, the estimated direct labor-hours were 21,940 hours and the total estimated manufacturing overhead was $530,948. At the end of the year, actual direct labor-hours for the year were 21,850 hours and the actual manufacturing overhead for the year was $530,948. Overhead at the end of the year was:

  • $2,178 overapplied

  • $2,178 underapplied

  • $2,228 underapplied

  • $2,228 overapplied

If manufacturing overhead is underapplied, then:

  • the amount of manufacturing overhead cost applied to Work in Process is less than the actual manufacturing overhead cost incurred.

  • the predetermined overhead rate is too high.

  • the Manufacturing Overhead account will have a credit balance at the end of the year.

  • actual manufacturing overhead cost is less than estimated manufacturing overhead cost.
Answer #1

Contribution Margin = 105,000*77% = $80,850

Less: Fixed expenses = 46,000

Net Operating Income = $34,850

2. · the total activity for the activity cost pool.

3. · total contribution margin equals total fixed expenses.

4.Predetermined rate = 530,948/21,940 = $24.2 per hour

Applied = 24.2*21,850 = $528,770

Actual Overheads = $530,948

Underapplied = $2,178

· $2,178 underapplied

· the amount of manufacturing overhead cost applied to Work in Process is less than the actual manufacturing overhead cost incurred.